OpenStudy (clara1223):

find f'(x) of f(x)=sin(x)+5cos(x)

OpenStudy (johnweldon1993):

\[\large \frac{d}{dx} sin(x) + 5cos(x) = \frac{d}{dx} sin(x) + \frac{d}{dx} 5cos(x)\]

OpenStudy (johnweldon1993):

Just a quick hint there*

OpenStudy (clara1223):

thanks, I'll see if i can figure it out from there

OpenStudy (johnweldon1993):

No problem! Just reply if you need help :)

OpenStudy (clara1223):

so would the answer be cos(x)-5sin(x)?